Question text

The motion of a wave traveling along a stretched string is described by the equation:

y(x,t)=(8.5cm)sin(5.2x−7.2t)

where x is in metres and tt is in seconds.

What is the minimum time it takes for a particle on the string to move from y = -8.5 cm to y = 8.5 cm?
Answer #1

Here angular frequency w = 7.2 s^-1

Minimum time = time taken to reach from maximum to minimum = T/2 = [2pi/w]/2 = pi/w = pi/7.2 = 0.436 s answer